# Online Continual Graph Learning

### v1

@@ III-A Problem formulation
Mini-batch acquisition. Within the introduced OCGL framework, and depending on the application, predictions may be requested from -hop neighbors with . However, due to the discussed neighborhood expansion problem, using all of these nodes does not allow operating within the computational efficiency constraints. This issue is similar to the problem of scaling static GNNs for large graphs, for which mini-batch training is required both for memory and efficiency reasons. Numerous approaches have been developed to address this problem, such as fixing a number of neighborhood sampled for aggregation [53, 54] or training on partitions of the graph [55]. The problem is also present in CGL, but it is not in contrast with the requirements of the setting, and is thus not discussed in the literature. In our context, the simplest solution to guarantee compliance with OCGL requirements appears to be providing a maximum number of neighboring nodes through any form of sub-sampling, thus guaranteeing an upper bound on the size of the computational graph for each batch. It follows that the OCGL model will not have direct access to the entire graph . For each node in a mini-batch, an OCGL model cannot leverage the full -layer ego-graph within . Instead, it will only received a sub-sampled .
@@
The way in which the sub-sampled -hop neighborhoods are obtained starting from the data generating process of the growing graph may depend on the specific application, but we can generally conceptualize two distinct systems, which are nonetheless equivalent regarding the learning process of the model. In one case, an up-to-date snapshot is assumed to be stored in a Past Information Store (PIS) system – distinct from an eventual memory buffer associated with predictive model – as in the more general lifelong-learning system setup [3]. We do not impose memory limitations on the PIS to allow graph growth, but we still require the training on each mini-batch to have bounded computational time and memory cost, and assume to only have access to the labels of nodes in the current batch. In the other case, mini-batches are directly received in the form of sub-sampled neighborhoods from an external data source. This could be the case when privacy concerns are involved, or in case of huge graphs where a local exploration is performed.
@@ IV Strategies for OCGL
Having defined the Online Continual Graph Learning setting, we consider and evaluate some popular CL techniques, most of which are agnostic with respect to the type of the input data. Some CGL learning strategies are easily adapted to the online setting, such as PDGNN [56] and SSM [57], while others are not directly applicable, such as ER-GNN [13], which stores representative nodes according to metrics computed offline on an entire graph snapshot, and thus we simplify it as described below. Similar efficiency considerations exclude many more recent CGL methods, which may require expensive steps and/or access to most of the graph: to name a few, CaT [58] requires access to the task subgraph for condensation, UGCL [59] uses a global structure distillation which requires to compute node embeddings for the entire graph, and a local structure distillation which still requires to compute the full embedding of buffer nodes and all their neighbors, and MSCGL [60] performs a costly neural architecture search for each task. Additionally, baselines that require expensive fine-tuning steps such as GDumb [61] are excluded, as they would violate the online setting. Overall, most strategies natively require task boundaries, and have been modified for the task-free setting as described below.
- 1.
2.
@@
3.
A-GEM. Averaged GEM [5] is a more efficient version of GEM [18], which ensures that the average loss for past tasks does not increase. It achieves this by projecting the gradient of the incoming batch in the orthogonal space of the gradient computed on samples from a memory buffer, if their scalar product is negative. We select buffer nodes with reservoir sampling [62].
- 4.
5.
@@ V Experimental Setup
In this section we introduce the specific experimental setup used, describing the construction of the node streams from benchmark datasets, and the details of model training and evaluation.111Code available at https://github.com/giovannidonghi/OCGL.
@@
Benchmarks. Six node classification graph datasets are used in our experiments: four homophilous multiclass datasets CoraFull [66], Arxiv [67], Reddit [53] and Amazon Computer [68], the heterophilous Roman Empire [69] and the binary classification dataset Elliptic [70]. The datasets are described in Appendix A. As is common in the OCL literature [6, 27], in order to position our experiments close to the rest of the CL literature, for all datasets except for Elliptic we devise a node stream derived from the class-incremental CL setting, which is considered as the most challenging one for catastrophic forgetting [21]. We divide the nodes in the graph into groups with fixed order consisting of 2 classes: this would be the sequence of two-class tasks in class-incremental learning (resulting in 35 tasks for CoraFull, 20 tasks for Arxiv and Reddit, 5 for Amazon Computer and 9 for Roman Empire). Then, we fix and ordering on the nodes of each task, and we stream the nodes accordingly. Therefore, the graph will gradually grow with mini-batches of nodes from two new classes at a time, which are processed in an online fashion. This allows us to consider metrics from the CL literature which require task boundaries, even though in our experiments the learning algorithm itself is task agnostic and simply adds a new output neuron when an instance of a new class is observed. For Elliptic, which only has two classes and has time information, we construct a time-incremental stream, that is according to the node real timestamps, and we divide the nodes into 10 tasks solely for the purposes of evaluation. For each dataset, we split the graph into 60% for training, 20% for validation and 20% for testing. A transductive setting is used: validation and test nodes are not used for loss computation, but they are still used for message passing.
@@
Performance assessment. We consider three metrics widely adopted in the literature: Average Performance (AP), Average Forgetting (AF) [18], and Average Anytime Performance (AAP) [71]. AAP is obtained by evaluating the model on the validation set after each training batch, which we refer to as anytime evaluation. For the highly unbalanced Elliptic dataset, we consider the F1 score as performance metric, while for all other datasets accuracy is used. More details are reported in Appendix C.
@@
Training details. In our experiments the backbone for all the CL strategies is the Graph Convolutional Network (GCN) [32], specifically a 2-layer GCN with a fixed hidden dimension of 256 units as done by Zhang et al. [46], with the exception of PDGNN which uses SGC [65] followed by a 2 layer MLP with 256 units. We use Adam optimizer [72] without weight decay, tuning the learning rate as a hyperparameter with the protocol defined below. We consider the batch size to be fixed, as it could depend on the real-world problem. For the smaller datasets CoraFull, Amazon Computer, and Roman Empire we consider batches of 10 nodes, while for the larger Arxiv, Reddit, and Elliptic we use 50. To evaluate the impact of batch size, in Appendix E we report experiments with larger batch sizes for each dataset. To address the neighborhood expansion problem raised in Section III-A, we perform neighborhood sampling. We fixed the number of neighbors to 10 for CoraFull, Amazon Computer, and Arxiv, and 25 for Reddit, as the latter is much more densely connected compared to the others, with an average degree of 492. As suggested by Aljundi et al. [73], multiple passes on the same mini-batch before passing to the next can be beneficial. We therefore considered as an additional tuned hyperparameter whether to perform multiple passes (5) on each mini-batch. We stress how multiple passes on an individual mini-batch are compliant with the constraints of the online setting, as they do not require storing past mini-batches beyond the current one and processing time is multiplied by a constant factor, the number of passes. As a baseline, we use a bare model that is simply fine-tuned on the incoming stream without applying any CL strategy. Additionally, we provide an upper bound in the form of a model that is jointly trained offline on the entire graph. For replay based methods, we consider different sizes of the memory buffer: for CoraFull, Amazon Computer, Roman Empire, and Elliptic we consider 1000 or 4000 nodes, and for the larger Arxiv and Reddit 1000, 4000 or 16000.
@@
Hyperparameter selection. Many works in the CL literature use a learning protocol that is akin to the classic machine learning setting, selecting hyperparameters by performing as many full passes over the task sequence as required by a grid search. This protocol violates stricter definitions of Lifelong Learning, where the stream is observed only once, and is indeed unrealistic for real applications where a model needs to quickly adapt to changes in data distribution. Chaudhry et al. [5] therefore proposed a more sensible hyperparameter selection protocol, which has now been used in several works [41, 6, 27] and that we use for our experiments. With this protocol, only the first few tasks are used for hyperparameter selection, allowing the model to perform multiple passes, with the same online setting, over them to select the hyperparameters that lead to the best performance on validation nodes. In our case we considered approximately 20% of the tasks for this validation, with the exception of Amazon Computer where it was set to 2 as there are only 5 total tasks. Hyperparameters are then selected based on the AP on the validation set at this validation boundary (details on the search space are reported in Appendix D).
@@ VI Results
In this section we discuss empirical results obtained on mini-batches with neighborhood sampling. Further results with larger batch sizes are reported in Appendix E, together with experiments with full neighborhood to assess the impact of sampling on model performance. A comparison of AP, AAP and AF for the considered CL methods on the six benchmark datasets is reported in Tables II-VII. The results for replay-based methods in the table are those obtained with the larger memory buffer sizes (4000 for CoraFull, Amazon Computer, Roman Empire, and Elliptic, 16000 for Arxiv and Reddit), while a comparison between different buffer sizes is shown in Figure 3 and in Appendix E. All experiments were repeated five times with different initializations, and the metrics are reported as mean and standard deviation across runs.
VI-A Homophilous dataset, class-incremental stream
We discuss here the results for the four homophilous datasets with a class-incremental stream, a setting that remains closer to the rest of the CGL literature [46, 57, 56].
Final Average Performance. Across all datasets, we observe that most of the compared strategies do not come close to the upper bound consisting of joint training. In general, the considered replay methods, specifically A-GEM, SSM-ER and PDGNN achieve higher final AP compared to the baseline and regularization methods, and often with a large margin. This can be expected, as rehearsal methods in CL are generally known to achieve most of the state-of-the-art results [14]; this holds true also for CGL [46]. Regularization methods struggle more than the replay strategies, as only on CoraFull EWC is significantly better than the bare baseline, while on other datasets their performances are closer to the lower bound.
Anytime Average Performance. Looking at Average Performance gives us an easy way to compare the performance after the entire learning process. However, since in the online setting we expect the model to be ready to make predictions at any time, it is arguably more useful to look at Average Anytime Performance as a metric of performance over time. For example, while on Reddit (Table IV) A-GEM records a higher final AP than PDGNN and SSM-ER, its AAP is considerably lower. Thus, in this case PDGNN and SSM-ER may be preferable, due to a more robust performance over time. For this purpose Figures 14-16 of Appendix F can give us some additional insights into the performance trends through the node stream. We note how it is natural and expected that accuracy tends to decrease with the batch index, as new classes are introduced and the classification task gets increasingly complex. In general, we observe that while the performance of regularization methods is almost monotonically decreasing, the performance of replay-based methods shows much higher variations.
Buffer size for replay methods. In Figure 3, we observe the results obtained by the replay methods with different buffer sizes. We note how SSM-ER appears to be the only method to consistently and greatly benefit from an increased buffer size. As for SSM-based buffers the number of examples is lower than the size of the buffer due to the storage of some neighboring nodes in the form of a sparsified subgraph, we can understand how with lower memory capacity this strategy may not be viable, potentially overfitting to a very small set of examples. PDGNN on the other hand shows a consistently high performance even with smaller buffer size, indicating a more efficient use of memory, as it is topology aware.
Strategy comparison. To better understand the different results in light of the choice of strategy, we report in Figure 4 a more detailed breakdown of accuracy by task for three representative methods on the Reddit dataset. With this comparison we clearly see how TWP (4a) struggles to retain past knowledge for most tasks, despite its regularization mechanism. Instead, A-GEM (4b) seems to achieve a better trade-off between stability and plasticity, maintaining the capacity to learn new tasks while generally preserving acquired knowledge. Finally, as discussed earlier PDGNN (4c) shows a more consistent performance through time, despite a lower final performance.
VI-B Heterophilous dataset, class-incremetal stream
To the best of our knowledge, the issue of heterophily in CGL has been addressed only by Zhao et al. [74]. However, their setting ignores inter-task edges, effectively making the graph less heterophilous. Since that requires unrealistic supervisory information on test nodes, we consider the Roman Empire dataset in our setting, showing results in Table VI. As can be expected, the results of all considered strategies are much lower than the upper bound. This can be the effect of two factors: first, the considered graph convolution may not be the best to deal with heteropily, and second, since nodes tend to connect with nodes of different classes, many connections are inter-task, and may thus not be observed when training on a certain class. This catastrophic drift in neighborhood composition is of course due to the class-incremental setting, which itself may not be realistic for heterophilous phenomena. While Roman Empire is not a real-world dataset, since it comes from words in a Wikipedia page, our results show that training GCNs in the OCGL setting poses significan challenges due to the drift in neighborhood composition.
| Method | AP% | AAPval% | AF% |
@@ VI-B Heterophilous dataset, class-incremetal stream
VI-C Time-incremental stream
Compared to the abrupt and challenging drifts of the class-incremental stream, on the Elliptic dataset we have a tamer distribution shift, determined by the node timestamps, which naturally reduces the usefulness of CL. This is reflected in the results of Table VII, where many methods exhibit positive forgetting, indicating backward transfer. We also observe how in terms of final AP no CL method performs better than simple finetuning in the OCGL setting, and replay methods seem to perform worse than regularization ones, likely due to overfitting to the memory buffer. On the other hand, considering AAP and looking at Figure 18 (Appendix F), we see how PDGNN maintains an overall better and more stable performance, only with a slight drop at the end of the stream. Thus, while in this case using CL methods may not be crucial as in the more pronounced drifts, using the right strategy can still provide benefits.
| Method | AP% | AAPval% | AF% |

### v2

@@ III-D Mini-batch acquisition
As discussed, the growth of node neighborhoods makes it infeasible to aggregate information from all -hop neighbors while maintaining a bounded computation. To satisfy the efficiency requirements, the model must therefore operate on a limited subset of nodes. This issue is similar to the problem of scaling static GNNs for large graphs, where mini-batch training is required both for memory and efficiency reasons. Numerous approaches have been developed, such as fixing a number of neighbors sampled for aggregation [52, 53] or training on partitions of the graph [54]. In our context, the simplest solution to guarantee compliance with OCGL requirements is to cap the number of neighboring nodes through any form of sub-sampling, thus guaranteeing an upper bound on the size of the computational graph for each batch. Thus, for any node , an OCGL model cannot leverage the full -layer ego-graph within . Instead, it will only receive a sub-sampled .
@@
The way in which the sub-sampled -hop neighborhoods are obtained starting from the data generating process of the growing graph may depend on the specific application, but we can generally conceptualize two distinct systems, which are nevertheless equivalent regarding the learning process of the model. In one case, an up-to-date snapshot is assumed to be stored in a Past Information Store (PIS) system – distinct from an eventual memory buffer associated with predictive model – as in the more general lifelong-learning system setup [55]. We do not impose memory limitations on the PIS to allow graph growth, but we still require the training on each mini-batch to have bounded computational time and memory cost, and assume to only have access to the labels of nodes in the current batch. In the other case, mini-batches are directly received in the form of sub-sampled neighborhoods from an external data source. This could be the case when privacy concerns are involved, or in case of huge graphs where a local exploration is performed.
@@ IV Strategies for OCGL
IV-A Adaptation of existing methods
Having defined the Online Continual Graph Learning setting, we consider and evaluate some popular CL techniques, most of which are agnostic with respect to the type of the input data. Some CGL learning strategies are easily adapted to the online setting, such as PDGNN [56] and SSM [57], while others are not directly applicable, such as ER-GNN [9], which stores representative nodes according to metrics computed offline on an entire graph snapshot, and thus we simplify it as described below. Similar efficiency considerations exclude many more recent CGL methods, which may require expensive steps and/or access to most of the graph: to name a few, CaT [58] requires access to the task subgraph for condensation, UGCL [59] uses a global structure distillation which requires to compute node embeddings for the entire graph, and a local structure distillation which still requires to compute the full embedding of buffer nodes and all their neighbors, and MSCGL [60] performs a costly neural architecture search for each task. Additionally, baselines that require expensive fine-tuning steps such as GDumb [61] are excluded, as they would violate the online setting. Several strategies natively require task boundaries, and have been modified for the task-free setting as described below.
A-GEM. Averaged GEM [3] is a more efficient version of GEM [14], which ensures that the average loss for past tasks does not increase. It achieves this by projecting the gradient of the incoming batch in the orthogonal space of the gradient computed on samples from a memory buffer, if their scalar product is negative. We select buffer nodes with reservoir sampling [62]
@@ IV-A Adaptation of existing methods
IV-B A new baseline method
The main purpose of this section is introducing a minimal, efficient and competitive method that can serves as reference for future OCGL research. Following the insights obtained from our results in Section VI, we introduce LINEAR (Lightweight Incremental NEighborhood Aggregation with Replay), a new baseline inspired by PDGNN and pushing its design philosophy to the limit with regard to simplicity. For each node, LINEAR first computes a feature vector by averaging features of its neighbors. A lightweight linear classifier is then applied to these aggregated features.
More specifically, LINEAR consists of a single SGC layer [64] that averages up to feature vectors sampled from the 1-hop neighborhoods, followed by a multinomial logistic regression classifier – no hidden layers, no nonlinearities, and no iterative message passing are employed; is a model hyperparameter to meet the requested computational budget as described in Section III-D. We couple this with an experience replay mechanism, storing directly the neighborhood-averaged features via reservoir sampling.
The computational complexity of LINEAR is significantly lower than standard GNN-based models. Assume for convenience fixed feature dimension and number of sampled neighbors throughout all layers, and a -class linear classifier on top. Due to message passing, an -layer GNN requires processing nodes to compute the embedding of a single target node. Each node transformation applies an linear map, and when this cost dominates both aggregation and classification. The overall complexity of predicting the class of a single node is therefore , reflecting the neighborhood expansion problem. In contrast, LINEAR requires only cost, coming from the neighbor aggregation and the final linear classifier, respectively.
Despite its simplicity, we show in Section VI that LINEAR exhibits strong empirical performance in the OCGL benchmarks, often surpassing GNN-based methods.
@@ V Experimental Setup
In this section, we introduce the specific experimental setup used, describing the construction of the node streams from benchmark datasets, and the details of model training and evaluation.
V-A Benchmarks
@@
Seven node classification graph datasets are used in our experiments: five homophilous multiclass datasets CoraFull [66], Arxiv [67], Reddit [52], Amazon Computer [68] and Products [67], the heterophilous Roman Empire [69] and the binary classification dataset Elliptic [70]. The datasets are described in Appendix A. As is common in the OCL literature [4, 24], in order to position our experiments close to the rest of the CL literature, for all datasets except for Elliptic we devise a node stream derived from the class-incremental CL setting, which is considered the most challenging one for catastrophic forgetting [17]. We divide the nodes in the graph into groups with fixed order consisting of 2 classes: this would be the sequence of two-class tasks in class-incremental learning (resulting in 35 tasks for CoraFull, 20 tasks for Arxiv and Reddit, 5 for Amazon Computer, 23 for Products and 9 for Roman Empire). Then, we fix an ordering on the nodes of each task, and we stream the nodes accordingly. Therefore, the graph will gradually grow with mini-batches of nodes from two new classes at a time, which are processed in an online fashion. This allows us to consider metrics from the CL literature which require task boundaries, even though in our experiments the learning algorithm itself is task agnostic and simply adds a new output neuron when an instance of a new class is observed. For Elliptic and Arxiv, that have time information available, we construct a time-incremental stream, that is according to the node real timestamps, and we divide the nodes into 10 tasks solely for the purposes of evaluation. For each dataset, we split the graph into 60% for training, 20% for validation and 20% for testing. A transductive setting is used: validation and test nodes are not used for loss computation, but they are still used for message passing.
V-B Performance assessment
@@
We consider three widely adopted metrics in the literature: Average Performance (AP), Average Forgetting (AF) [14], and Average Anytime Performance (AAP) [71]. AAP is obtained by evaluating the model after each training batch, which we refer to as anytime evaluation, and thus allows us to understand model performance over time. For the highly unbalanced Elliptic dataset, we consider the F1 score as performance metric, while for all other datasets accuracy is used. More details are reported in Appendix B.
V-C Training details
@@
In our experiments the backbone for all CL strategies is the Graph Convolutional Network (GCN) [29], specifically a 2-layer GCN with a fixed hidden dimension of 256 units as done by Zhang et al. [45], with the exception of PDGNN which uses 2 layers of SGC [64] followed by a 2 layer MLP with 256 units, and LINEAR which simply averages features in the immediate neighborhood and uses a linear classifier on top. We use Adam optimizer [72] without weight decay, tuning the learning rate as a hyperparameter with the protocol defined below. We consider the batch size to be fixed, as it could depend on the real-world problem. For the smaller datasets CoraFull, Amazon Computer, and Roman Empire we consider batches of 10 nodes, while for the larger Arxiv, Reddit, and Elliptic we use 50, and for the much larger Products we use 250 nodes. To address the neighborhood expansion problem raised in Section III, we perform neighborhood sampling, fixing the number of neighbors to 10. As suggested by Aljundi et al. [73], multiple passes on the same mini-batch before passing to the next can be beneficial. We therefore considered as an additional tuned hyperparameter whether to perform multiple passes (5) on each mini-batch. We stress how multiple passes on an individual mini-batch are compliant with the constraints of the online setting, as they do not require storing past mini-batches beyond the current one and processing time is multiplied by a constant factor, the number of passes. As a baseline, we use a bare model that is simply fine-tuned on the incoming stream without applying any CL strategy. Additionally, we provide an upper bound in the form of a model that is jointly trained offline on the entire graph. For replay based methods, we consider different sizes of the memory buffer: 1%, 2% and 4% of the nodes, except for the much larger Products where we use 0.1%, 0.2% and 0.4%, to avoid having an unrealistically large buffer of several tens of thousands of nodes.
V-D Hyperparameter selection
@@
Many works in the CL literature use a learning protocol that is akin to the classic machine learning setting, selecting hyperparameters by performing as many full passes over the task sequence as required by a grid search. This protocol violates stricter definitions of Lifelong Learning, where the stream is observed only once, and is indeed unrealistic for real applications where a model needs to quickly adapt to changes in data distribution. Chaudhry et al. [3] therefore proposed a more sensible hyperparameter selection protocol, which has now been used in several works [40, 4, 24] and that we use for our experiments. With this protocol, only the first few tasks are used for hyperparameter selection, allowing the model to perform multiple passes, with the same online setting, over them to select the hyperparameters that lead to the best performance on validation nodes. In our case, we considered approximately 20% of the tasks for this validation, with the exception of Amazon Computer where it was set to 2 as there are only 5 total tasks. Hyperparameters are then selected based on AP on the validation set at this validation boundary (details on the hyperparameters are reported in Appendix C).
@@ VI Results
| Method | AAP% | AP% | AF% |
| A-GEM | |||
| ER | |||
@@
In this section, we discuss empirical results of the experiments described in Section V. A comparison of AAP, AP and AF for the considered CL methods on the various benchmarks is reported in Tables II-VIII. The results for replay-based methods in the tables are those obtained with the larger memory buffer sizes (4% of nodes, or 0.4% on Products), while a comparison between different buffer sizes is shown in Figure 5 and in Appendix E. All experiments were repeated five times with different initializations, and the metrics are reported as mean and standard deviation across runs. In the tables, the best results are highlighted in bold, the second-best underlined, and the third with dashed underline.
VI-A Overall results
Across benchmarks, the results confirm the difficulty of the OCGL setting, especially for a class-incremental stream: none of the considered strategies approaches the upper bound of AP provided by joint offline training, except for the proposed LINEAR on Amazon Computer. In general, the considered replay methods, specifically SSM-ER, PDGNN and LINEAR achieve higher performance compared to the baseline and regularization methods, and often with a large margin. This can be expected, as rehearsal methods in CL are generally known to achieve most of the state-of-the-art results [10]; this holds true also for CGL [45]. Regularization methods struggle more than the replay strategies, as only in few cases, such as with MAS on CoraFull, they score significantly better than the bare baseline, while their performances are often closer to the lower bound of simple fine-tuning.
VI-B Anytime evaluation
Looking at Average Performance gives us an easy way to compare the performance after the entire learning process. However, since in the online setting we expect the model to be ready to make predictions at any time, Average Anytime Performance is a more useful metric of performance over time. Although overall AP and AAP rankings agree, there are cases in which they differ. Figure 3 can give us some additional insights into the performance trends through the node stream. For example, on CoraFull we have that LINEAR scores a lower final AP compared to PDGNN, despite having a higher AAP. Yet, looking at Figure 3a we see that this is due to a drop in performance just at the end of the stream, while for the rest of the time LINEAR stays consistently above PDGNN. Or in Figure 3b, for the Reddit dataset, we see how while LINEAR overall closely follows PDGNN, it suffers from significantly more contained drops at task boundaries. Thus, these cases confirm AAP as more suitable metric than AP, as the behavior of LINEAR is preferable for anytime evaluation.
VI-C Strategy comparison
To better understand the different results in light of the choice of strategy, we report in Figure 4 a more detailed breakdown of accuracy by task for three representative methods on the CoraFull dataset. With this comparison we clearly see how MAS (4a) retains past knowledge for some tasks thanks to its regularization, but struggles to learn new information. Instead, PDGNN (4b) and LINEAR (4c) strike a better balance between stability and plasticity. They obtain similar results, but as observed above LINEAR appears more robust to sudden performance drops compared to PDGNN.
VI-D Time-incremental stream
Compared to the abrupt and challenging drifts of the class-incremental stream, with a time-incremental stream, determined by the node timestamps, we have a tamer distribution shift which naturally reduces the usefulness of CL. This is reflected in the results of Tables VII-VIII, where many methods exhibit positive forgetting, indicating some backward transfer. Although CL offers limited improvement in these conditions, considering AAP and looking at appendix Figure 12 (Appendix E), we see how PDGNN maintains an overall better and more stable performance, only with a slight drop at the end of the stream. Thus, while on these two dataset using CL methods may not be crucial as in cases with more pronounced drifts, using the right strategy can still provide benefits.
| Method | AAP% | AP% | AF% |
| A-GEM | |||
| ER | |||